17. A 0.100 m solution of which one of the following solutes will have the lowest

vapour pressure?

A. KClO4
B. Ca(ClO4)2

C. Na3PO4

D. C6H12O6

E. NaCl
1
Expert's answer
2021-08-26T07:16:40-0400

Solution


KClO4 = 1K+ and 1ClO4: i=2

2"\times 0.1= 0.2"

Ca(Cl04)= 1Ca2+ and 2ClO4- : i=3

"3\times 0.1= 0.3"

Al(ClO4)3= 1Al3+ and 3ClO4-: i=4

"4\times 0.1= 0.4"

Sucrose i= 1

"1\times0.1= 0.1"

1Na+ and 1Cl- :i=2

"2\times0.1 =0.2"


Therefore Al(ClO4)3 has the lowest vapour pressure of 0.4